OpenStudy (anonymous):

5/7-3/7 7/8-3/8 9/11-7/11

OpenStudy (anonymous):

HELP MEEEEEEEEEEEEEEE

OpenStudy (diyadiya):

5/7-3/7=2/7 7/8-3/8=4/8 =1/2 9/11-7/11=2/11